OpenStudy (jenny.andrade):

Express this quadratic equation as the product of two binomials. y = 4x^2 - 28x - 32

OpenStudy (ikram002p):

factorize it

OpenStudy (ikram002p):

y = 4(x^2 - 7x - 8 )

OpenStudy (ikram002p):

y=4(x+1)(x-8)

OpenStudy (jenny.andrade):

thankyou how do you do this one? Find the roots of the equation y = 4x^2 - 20x - 24 .

OpenStudy (ikram002p):

first factorize it can you do that ?

OpenStudy (jenny.andrade):

No i don't k now how?

OpenStudy (jenny.andrade):

know*

OpenStudy (ikram002p):

\( y = 4x^2 - 20x - 24 . =4(x^2-5x-6)=4(x-6)(x+1)\) nw 4(x-6)(x+1)=0 we have two cases x-6=0 \(\rightarrow \) x=6 x+1=0 \(\rightarrow \) x=-1

OpenStudy (jenny.andrade):

So (-1.0) and (6,0) ?

OpenStudy (ikram002p):

yep ^

OpenStudy (jenny.andrade):

Okay i also need help on this one Find c such that x^2 - 24x + c is a perfect square trinomial.

OpenStudy (ikram002p):

well i feel tired nw post it in new thread im sure someone else will help good luck ^

OpenStudy (jenny.andrade):

OKay thankyou for your help(: